# HalluVault: A Novel Logic Programming-aided Metamorphic Testing Framework for Detecting Fact-Conflicting Hallucinations in Large Language Models

### v1

@@ 5.3. RQ2: FCH Categorization and Analysis
We categorize the hallucination responses in more detail and focus primarily on two types of hallucination: error knowledge response, error inference response, and contradictory response. Note that we consider refusal to respond such as ‘I don’t know’ due to the lack of relevant knowledge as adhering to LLMs’ honesty and truthfulness principles. Therefore, we categorize refusal to respond as a normal response. To ensure fair and unbiased categorization, 100 hallucination-related responses were randomly selected and independently categorized by three co-authors, who then discussed the results to reach a consensus categorization.
@@
Here we provide the distribution of the hallucination categorization results, as demonstrated in Figure 7. There is partial overlap between these two types of hallucinations because incorrect reasoning processes may also involve erroneous knowledge. Among these issues, there are several contradictory answers primarily arising from inconsistency between incorrect reasoning processes and correct answers; thus, it exists in these two types of errors. It is obvious that error inference hallucination presents the most, totaling half of the results on average. This indicates that the primary cause of FCH issues in logical reasoning is the insufficient reasoning capability of LLMs. Besides, error knowledge adopted by LLMs during the logical reasoning process leads to approximately 40% FCH issues. The overlaps account for about 8%-20% at the hallucination ratio, which indicates there are entities where LLMs have learned entirely erroneous relevant information, necessitating the employment of certain measures for correction.
@@ 5.3.3. Case Study
Error Inference Hallucination. One of the most common types of logical reasoning leading to error inference hallucination is temporal attribute reasoning, proven to be a category of reasoning task that performs poorly on LLMs (Qiu et al., 2023). Experiments on time-related reasoning tasks are comprehensively conducted and unsatisfactory performance of LLMs are observed.
@@
Error Knowledge Hallucination. Figure 9 demonstrates a classic example of LLM hallucination caused by using error knowledge for logical reasoning. General Dmitry Karbyshev (1880-1945) was a Russian Imperial Army soldier who served in several wars during World War I (1914-1918) and II (1939-1945), and Louis Bernacchi (1876-1942) was an Australian physicist and astronomer who served in the Royal Naval Volunteer Reserve during World War I. Thus, the ground truth answer to this question should be ‘Yes’. However, when testing with Llama2-7B-chat-hf, an inspiring observation is that when LLMs encounter unfamiliar knowledge, they do not adhere to the honesty principle; instead, they fabricate knowledge and its sources. We subsequently employ an RAG-based scheme to reintroduce relevant knowledge to the model, leading to the restoration of normal responses.
@@
We further conduct an out-of-distribution (OOD) knowledge experiment to figure out the cause of error knowledge hallucination. OOD is another factor that could cause FCH issues (Zhang et al., 2023). We design contextual reasoning utilizing recent sporting events and natural disasters from Wikipedia since June 2023, which is considered unutilized information in LLMs’ training data based on their up-to-date introductions. We construct a series of test cases containing contextual descriptions of recent events using HalluVault, observing whether LLMs can be guided to respond to OOD knowledge and trigger FCH.
@@
Figure 10 is a typical case of OOD contexts leading to error knowledge hallucination. In the initial test of GPT-3.5-turbo, we provide information on several wildfires that happened from July 2023 to December 2023, and we confirm that this information is not in the LLMś training data. The LLM subsequently indicates that it has acquired this knowledge through this interactive process. However, a turning point emerges when we use test cases designed by HalluVault in the second test. Despite our questions based on preliminary factual knowledge provided, the LLM still confidently responds with a wrong answer.
@@
We analyze several potential causes for this situation. One possibility is that LLMs store incorrect knowledge in the first turn because what we provided was merely a list of events, rather than a list of events in their order of occurrence. In short, the normal reasoning process involves defining the earliest occurring events only after knowing the times of all events. However, the LLM opts to judge based on the order we provide event knowledge, which is contrary to facts. Another potential is that when LLMs encounter OOD knowledge if they do not strictly adhere to the principle of honesty by stating I do not know…, they tend to complete and analyze the response based on error knowledge in their existing knowledge bases. Nevertheless, such responses are likely to induce hallucinations.
5.4. RQ3: Ablation Study
@@
We conduct an ablation study to investigate the capacity of each inference rule so that they can be distinctly used to uncover anomalies. The four types of rules illustrated in Section 4.2 are separately applied to generate Q&A pairs. The symmetric reasoning rule is primarily utilized within the composition reasoning rule and does not introduce new knowledge on its own. Therefore, we did not include the symmetric reasoning rule as a separate condition in our ablation study. For better visualization and understanding, we present the distribution of hallucination-related responses discovered with diverse rule-generated questions by HalluVault in Figure 11. The figure illustrates which type of rule can trigger the most hallucination responses for different LLMs and different domains of knowledge. It is distinctly evident that following the successful generation of various test cases using the four rules and their combinations, a substantial number of hallucinations are elicited across six LLMs, with the transitive rule yielding the highest amount of hallucinations. Following closely behind are the test cases generated using composition rules, which have triggered a significant number of FCHs in both the people and history domains.
@@ 6.1. Threats to Validity
Limited Coverage of Knowledge Databases. Our research predominantly employs data from the Wikipedia database to generate test cases using HalluVault. However, it is important to note that HalluVault is not limited to this specific database. Its design allows for easy extension and adaptation to various other knowledge bases, illuminating its versatility and applicability.
@@ 6.2. Mitigation
We involve two model editing algorithms, i.e., ROME (Meng et al., 2022a) and MEMIT (Meng et al., 2022b), to integrate new knowledge derived from reasoning into open-source LLMs, aiming to alleviate FCH issues. We apply FastEdit (hiyouga, 2023) and EasyEdit (Wang et al., 2023) for more speedy implementation. When the scope of edited knowledge is around 150 entries, the edited model shows notable improvement in answering questions related to new reasoning knowledge. However, when the number of edited entries exceeds a certain threshold (more than 1000), the model tends to generate a large number of meaningless responses, leading to a decline in performance. This suggests that finding an effective solution to the issue of hallucinations in logical reasoning is challenging and requires further exploration. Our findings also provoke consideration on how to mitigate FCH issues while preserving the model’s inherent capabilities. Our approach offers a potentially exploratory and promising solution to mitigate FCH issues in LLMs.
@@ 6.3. Takeaway Messages
Towards In-depth Understanding of LLM Hallucination. From the insights derived in this work, it is important to further explore techniques to understand the deep-rooted causes of hallucinations LLMs through white-box methods. A promising direction is to enhance and augment the logical reasoning capabilities of LLMs to reduce hallucination issues.
@@ 7. Related Work
LLM security and its testing has been studied by numerous work in the literature (Wang et al., 2024; Li et al., 2024; Deng et al., 2024; Liu et al., 2023). In this section, we briefly introduce some related or prior works on three topics listed below.
@@ 7.1. Evaluating Hallucination in Large Language Models
TruthfulQA (Lin et al., 2022) is the most classic dataset for assessing whether language models generate truthful answers to questions. It tests whether the models learn incorrect answers during the generation process due to emulating human text. Another dataset HaluEval (Li et al., 2023a) samples 10K instances from the training sets of HotpotQA (Yang et al., 2018), OpenDialKG (Moon et al., 2019), and CNN/DailyMail (See et al., 2017), and utilizes LLMs to generate hallucination-corresponding samples by setting tasks and employing specific sampling strategies, which is primarily aimed at question-answering tasks and text summarization tasks. KoLA (Yu et al., 2023) tests the hallucination issues of LLMs in the domain of knowledge graphs and introduces tasks based on 19 focal entities, concepts, and events. It assesses the capacity of large language models (LLMs) to handle structured knowledge across four levels: memory, understanding, application, and creation. This aims to test the hallucination phenomena of LLMs in the domain of knowledge graphs. From the perspective of long context, BAMBOO (Dong et al., 2023) and FActScore (Min et al., 2023) both target the long text generation capabilities of large language models, assessing their performance in extended context scenarios through factual verification. Additionally, there are assessments of large language models for hallucination issues in specific domains such as healthcare and finance (Umapathi et al., 2023; Kang and Liu, 2023).
@@ 7.2. Mitigating Hallucination in Large Language Models
Considerable work (Lightman et al., 2023; Varshney et al., 2023; Gou et al., 2023; Vu et al., 2023) involves utilizing external knowledge retrieval or automated feedback adjustments to make text responses from large language models more controllable and reliable. Similar approaches are proposed for multimodal hallucination mitigation such as Woodpecker (Yin et al., 2023), which extracts key concepts to generate questions and knowledge assertions for hallucination diagnosis and mitigation. Another thread involves using fine-tuning techniques to mitigate model hallucinations. AlpaGasus (Chen et al., 2023), Elaraby et al. (Elaraby et al., 2023) and Tian et al. (Tian et al., 2023) apply fine-tuning techniques on high-quality data for better effectiveness and factuality. Besides, the findings of Elaraby et al. (Elaraby et al., 2023) reveal that the knowledge injection technique enhances the performance of less robust LLMs. Additionally, an increasing number of researchers are turning towards studying white-box repairing methods for open-source large language models. The evidence presented in the discourse by Azaria et al. (Azaria and Mitchell, 2023) suggests that the internal states of Large Language Models can be utilized to discern the veracity of statements, thereby elucidating the underlying causes of factual hallucinations in LLMs. Studies like IIT (Li et al., 2023b) and Repr (Zou et al., 2023) endeavor to alleviate hallucination issues by delving into LLMs’ deep-layer information through the analysis of internal model states. This approach not only augments the interpretability of large language models but is also regarded as a vital research direction for the future of explainable and trustworthy AI.
7.3. Comparison with Existing Works
We qualitatively compare HalluVault with the state-of-the-art FCH evaluation approaches and existing natural language reasoning benchmarks to illustrate the advantages of HalluVault. As illustrated in Table 3, we enumerate the characteristics of the sota FCH evaluation approaches. To assess FCH in LLM responses, existing approaches uniformly opt for a wiki-related knowledge base as the foundation for constructing ground truth facts. Their main distinction from HalluVault lies in the manner of task construction and the metrics employed to measure hallucinations.
Task Construction Methods. Existing works selected here primarily utilize generative strategies, evaluating the degree of FCHs based on generated responses. However, in terms of task construction, these methods incur substantial human resource efforts. Apart from the KoLA-KM, KA (Yu et al., 2023), which is essentially a collection of existing Q&A datasets, both TruthfulQA (Lin et al., 2022) and HaluEval (Li et al., 2023a) rely on human annotations to construct Q&A pairs. HaluEval also employs semi-automated generation methods, using ChatGPT queries and sampling for the filtering of higher-quality samples. HalluVault, on the other hand, utilizes Prolog-assisted automatic inference to derive new knowledge triples and generate templates for new questions, achieving maximum automation of construction while ensuring the complexity of the questions.
Response Evaluation Metrics. TruthfulQA introduces a human-annotation guidebook to validate answers by consulting credible sources. Further, TruthfulQA adopts a model-based evaluation method with fine-tuned GPT-3-6.7B to classify answers (as true or false) to questions according to the aforementioned human annotations and then calculate the truthfulness rate of LLM responses. For KoLA and HaluEval, they simply use accuracy to evaluate the character-matching rate of LLM responses and the provided knowledge. Thus, HalluVault considers the structural similarity of LLM responses with original knowledge triples and the reasoning process, offering superiority over those simple evaluation metrics.
Dataset | Fact Source | Construction Method | Test Oracle |
TruthfulQA | Wikipedia pages & websites | Human annotations | Truthfulness Rate |
KoLA-KM, KA | Wikidata5M & websites | Existing datasets consolidation | Standardized score (F1) |
HaluEval-QA | Wikipedia | Human annotations & ChatGPT query | Accuracy |
HalluVault-Dataset | Wikidata triples | Prolog-aided reasoning & template-based generation | Structural Similarity |
As listed in Table 4, we provide several benchmarks aided for natural language reasoning. Existing reasoning benchmarks lean more towards logical predicate-formatted inputs and outputs, lacking natural language-formatted questions, thus limiting their suitability for testing with LLMs.
Benchmark | Size | Reasoning Type | Data Source | Task | Automation |
DEER | 1.2k | Inductive reasoning | Wikipedia | Rule Generation | ✗ |
HalluVault | Scalable | Deductive reasoning | Wikidata | Question Answering | ✓ |

### v2

@@ 5.3. RQ2: FCH Categorization and Analysis
We categorize the hallucination responses in more detail and focus primarily on two types of hallucination: error knowledge response and error inference response. Note that we consider refusal to respond such as ‘I don’t know’ due to the lack of relevant knowledge as adhering to LLMs’ honesty and truthfulness principles. Therefore, we categorize refusal to respond as a normal response. To ensure fair and unbiased categorization, 100 hallucination-related responses were randomly selected and independently categorized by three co-authors, who then discussed the results to reach a consensus categorization.
@@
Here we provide the distribution of the hallucination categorization results, as demonstrated in Figure 7. There is partial overlap between these two types of hallucinations because incorrect reasoning processes may also involve erroneous knowledge. Among these issues, there are several contradictory answers primarily arising from inconsistency between incorrect reasoning processes and correct answers; thus, it exists in these two types of errors. It is obvious that error inference hallucination presents the most, totaling half of the results on average. This indicates that the primary cause of FCH issues in logical reasoning is the insufficient reasoning capability of LLMs. Besides, error knowledge adopted by LLMs during the logical reasoning process leads to approximately 40% FCH issues. The overlaps account for about 8%-21% at the hallucination ratio, which indicates there are entities where LLMs have learned entirely erroneous relevant information, necessitating the employment of measures for correction.
@@ 5.3.3. Case Study
Error Inference Hallucination. One of the most common types of logical reasoning leading to error inference hallucination is temporal attribute reasoning, proven to be a category of reasoning task that performs poorly on LLMs (Qiu et al., 2024). Experiments on time-related reasoning tasks are comprehensively conducted and unsatisfactory performance of LLMs are observed.
@@
Error Knowledge Hallucination. Figure 9 demonstrates a classic example of LLM hallucination caused by using error knowledge for logical reasoning. General Dmitry Karbyshev (1880-1945) was a Russian Imperial Army soldier who served in several wars during World War I (1914-1918) and II (1939-1945), and Louis Bernacchi (1876-1942) was an Australian physicist and astronomer who served in the Royal Naval Volunteer Reserve during World War I. Thus, the ground truth answer to this question should be ‘Yes’. However, when testing with Llama2-7B-chat-hf, an inspiring observation is that when LLMs encounter unfamiliar knowledge, they do not adhere to the honesty principle; instead, they fabricate knowledge and its sources. We subsequently employ an RAG-based scheme to reintroduce relevant knowledge, leading to the restoration of normal responses.
@@
We further conduct an out-of-distribution (OOD) knowledge experiment to figure out the cause of error knowledge hallucination. OOD is another factor that could cause FCH issues (Zhang et al., 2023). We design contextual reasoning utilizing recent sporting events and natural disasters from Wikipedia since June 2023, which is considered unutilized information in LLMs’ training data based on their up-to-date introductions. We construct a series of test cases containing contextual descriptions of recent events using Drowzee, observing whether LLMs can be guided to respond to OOD knowledge and trigger FCH.
@@
Figure 10 is a typical case of OOD contexts leading to error knowledge hallucination. In the initial test of GPT-3.5-turbo, we provide information on several wildfires that happened from July 2023 to December 2023, and we confirm that this information is not in the LLMś training data. The LLM subsequently indicates that it has acquired this knowledge through this interactive process. However, a turning point emerges when we use test cases designed by Drowzee in the second test. Despite our questions based on preliminary factual knowledge provided, the LLM still confidently responds with a wrong answer.
@@
We analyze several potential causes for this situation. One possibility is that LLMs store incorrect knowledge in the first turn because what we provided was merely a list of events, rather than a list of events in their order of occurrence. In short, the normal reasoning process involves defining the earliest occurring events only after knowing the times of all events. However, the LLM opts to judge based on the order we provide event knowledge, which is contrary to facts. Another potential is that when LLMs encounter OOD knowledge if they do not strictly adhere to the principle of honesty by stating I do not know…, they tend to complete the response based on error knowledge in their existing knowledge bases. Nevertheless, such responses are likely to induce hallucinations.
5.4. RQ3: Comparison with Existing Works
5.4.1. Qualitative Analysis.
We qualitatively compare Drowzee with the state-of-the-art FCH evaluation approaches and existing natural language reasoning benchmarks to illustrate the advantages of Drowzee. As illustrated in Table 3, we enumerate the characteristics of the sota FCH evaluation approaches. Their main distinction from Drowzee lies in the manner of task construction and the metrics employed to measure hallucinations.
Task Construction Methods. Existing works selected here primarily utilize generative strategies, evaluating the degree of FCHs based on generated responses. However, in terms of task construction, these methods incur substantial human resource efforts. Apart from the KoLA-KM, KA (Yu et al., 2024), which is essentially a collection of existing Q&A datasets, both TruthfulQA (Lin et al., 2022) and HaluEval (Li et al., 2023a) rely on human annotations to construct Q&A pairs. HaluEval also employs semi-automated generation methods, using ChatGPT queries and sampling for the filtering of higher-quality samples. Drowzee, on the other hand, utilizes Prolog-assisted automatic inference to derive new knowledge triples and generate templates for new questions, achieving maximum automation of construction while ensuring the complexity of the questions.
Response Evaluation Metrics. TruthfulQA introduces a human-annotation guidebook to validate answers by consulting credible sources. Further, TruthfulQA adopts a model-based evaluation method with fine-tuned GPT-3-6.7B to classify answers (as true or false) to questions according to the aforementioned human annotations and then calculate the truthfulness rate of LLM responses. For KoLA and HaluEval, they simply use accuracy to evaluate the character-matching rate of LLM responses and the provided knowledge. FActScore (Min et al., 2023) is a method for evaluating the factuality of long texts generated by language models. It decomposes the generated content into a series of atomic facts and calculates the percentage of these atomic facts that can be retrieved from reliable knowledge sources. Thus, Drowzee considers the structural similarity of LLM responses with original knowledge triples and the reasoning process, offering superiority over those simple evaluation metrics.
For natural language reasoning scenarios, we provide several benchmarks as listed in Table 4. FOLIO (Han et al., 2022) is a natural language reasoning dataset annotated with first-order logic (FOL) by human experts, primarily used to test the deductive reasoning capabilities of generative language models. DEER (Yang et al., 2024b), on the other hand, focuses on the inductive reasoning paradigm, where natural language rules are induced from natural language facts, providing rule-fact pairs to test the inductive reasoning abilities of language models. Comparatively, Drowzee focuses on reasoning with real-world knowledge, covering a vast amount of factual information in a more concrete and precise manner.
Dataset | Fact Source | Construction Method | Test Oracle | Result (%) |
TruthfulQA | Wikipedia pages & websites | Human annotations | Truthfulness Rate | 89 |
KoLA-KM, KA | Wikidata5M & websites | Existing datasets consolidation | Standardized Score (F1) | 82 |
HaluEval-QA | Wikipedia | Human annotations & ChatGPT query | String Matching | 85 |
FActScore | Wikipedia | — | Atomic Fact & Retrieval | 97 |
Drowzee-Dataset | Wikidata triples | Prolog-aided reasoning & generation | Semantic Similarity | 100 |
Benchmark | Size | Reasoning Type | Data Source | Task | Automation |
DEER | 1.2k | Inductive reasoning | Wikipedia | Rule Generation | ✗ |
Drowzee | Scalable | Deductive reasoning | Wikidata | Question Answering | ✓ |
5.4.2. Small-scale Quantitative Analysis.
To evaluate the detection accuracy of Drowzee in comparison with existing methods, we conduct a small-scale quantitative analysis using a set of 100 test cases that are already manually verified. The success rates of this comparison are summarized in the last column of Table 3.
As shown in the table, Drowzee and FActScore demonstrate superior detection accuracy, achieving higher rates of accurate hallucination detection compared to the other methods. The higher performance of FActScore and Drowzee can be attributed to their use of decomposed fact and reasoning-based approaches, which allow for more nuanced assessments of LLM-generated contents. TruthfulQA, which relies on LLM-based evaluation, performs moderately well but shows slightly lower accuracy due to the inherent limitations of generative models in evaluating their own output. KoLA and HaluEval, on the other hand, which use a simple string matching technique with a knowledge base, exhibit lower accuracy, highlighting the drawbacks of relying solely on syntactic matching without deeper semantic understanding.
This quantitative analysis further underscores the advantages of Drowzee in providing a more reliable and scalable method for FCH detection in large language models.
5.5. RQ4: Ablation Study
@@
We conduct an ablation study to investigate the capacity of each inference rule so that they can be distinctly used to uncover anomalies. The four types of rules illustrated in Section 4.2 are separately applied to generate Q&A pairs. The symmetric reasoning rule is primarily utilized within the composite reasoning rule and does not introduce new knowledge on its own. Therefore, we did not include the symmetric reasoning rule as a separate condition in our ablation study. For better visualization and understanding, we present the distribution of hallucination-related responses discovered with diverse rule-generated questions by Drowzee in Figure 11. The figure illustrates which type of rule can trigger the most hallucination responses for different LLMs and different domains of knowledge. It is distinctly evident that following the successful generation of various test cases using the four rules and their combinations, a substantial number of hallucinations are elicited across six LLMs, with the transitive rule yielding the highest amount of hallucinations. Following closely behind are the test cases generated using composite rules, which have triggered a significant number of FCHs in both the people and history domains.
@@ 6.1. Threats to Validity
Limited Coverage of Knowledge Databases. Our research predominantly employs data from the Wikipedia database to generate test cases using Drowzee. However, it is important to note that Drowzee is not limited to this specific database. Its design allows for easy extension and adaptation to various other knowledge bases, illuminating its versatility and applicability.
@@ 6.2. Mitigation
We involve two model editing algorithms, i.e., ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023), to integrate new knowledge derived from reasoning into open-source LLMs, aiming to alleviate FCH issues. We apply FastEdit (hiyouga, 2023) and EasyEdit (Wang et al., 2024) for more speedy implementation. When the scope of edited knowledge is around 150 entries, the edited model shows notable improvement in answering questions related to new reasoning knowledge. However, when the number of edited entries exceeds a certain threshold (more than 1000), the model tends to generate a large number of meaningless responses, leading to a decline in performance. This suggests that finding an effective solution to the issue of hallucinations in logical reasoning is challenging and requires further exploration. Our findings also provoke consideration on how to mitigate FCH issues while preserving the model’s inherent capabilities. Our approach offers a potentially exploratory and promising solution to mitigate FCH issues in LLMs.
@@ 6.3. Takeaway Messages
Towards In-depth Understanding of LLM Hallucination. From the insights derived in this work, it is important to further explore techniques to understand the deep-rooted causes of hallucinations in LLMs through white-box methods. A promising direction is to enhance and augment the logical reasoning capabilities of LLMs to reduce hallucination issues.
@@ 7.1. Evaluating Hallucination in Large Language Models
TruthfulQA (Lin et al., 2022) is the most classic dataset for assessing whether language models generate truthful answers to questions. It tests whether the models learn incorrect answers during the generation process due to emulating human text. Another dataset HaluEval (Li et al., 2023a) samples 10K instances from the training sets of HotpotQA (Yang et al., 2018), OpenDialKG (Moon et al., 2019), and CNN/DailyMail (See et al., 2017), and utilizes LLMs to generate hallucination-corresponding samples by setting tasks and employing specific sampling strategies, which is primarily aimed at question-answering tasks and text summarization tasks. KoLA (Yu et al., 2024) tests the hallucination issues of LLMs in the domain of knowledge graphs and introduces tasks based on 19 focal entities, concepts, and events. It assesses the capacity of large language models (LLMs) to handle structured knowledge across four levels: memory, understanding, application, and creation. This aims to test the hallucination phenomena of LLMs in the domain of knowledge graphs. From the perspective of long context, BAMBOO (Dong et al., 2024) and FActScore (Min et al., 2023) both target the long text generation capabilities of large language models, assessing their performance in extended context scenarios through factual verification. Additionally, there are assessments of large language models for hallucination issues in specific domains such as healthcare and finance (Pal et al., 2023; Kang and Liu, 2023).
@@ 7.2. Mitigating Hallucination in Large Language Models
Considerable work (Lightman et al., 2024; Varshney et al., 2023; Gou et al., 2024; Vu et al., 2024) involves utilizing external knowledge retrieval or automated feedback adjustments to make text responses from large language models more controllable and reliable. Similar approaches are proposed for multimodal hallucination mitigation such as Woodpecker (Yin et al., 2023), which extracts key concepts to generate questions and knowledge assertions for hallucination diagnosis and mitigation. Another thread involves using fine-tuning techniques to mitigate model hallucinations. AlpaGasus (Chen et al., 2024), Elaraby et al. (Elaraby et al., 2023) and Tian et al. (Tian et al., 2024) apply fine-tuning techniques on high-quality data for better effectiveness and factuality. Besides, the findings of Elaraby et al. (Elaraby et al., 2023) reveal that the knowledge injection technique enhances the performance of less robust LLMs. Additionally, an increasing number of researchers are turning towards studying white-box repairing methods for open-source large language models. The evidence presented in the discourse by Azaria et al. (Azaria and Mitchell, 2023) suggests that the internal states of Large Language Models can be utilized to discern the veracity of statements, thereby elucidating the underlying causes of factual hallucinations in LLMs. Studies like IIT (Li et al., 2023b) and Repr (Zou et al., 2023) endeavor to alleviate hallucination issues by delving into LLMs’ deep-layer information through the analysis of internal model states. This approach not only augments the interpretability of large language models but is also regarded as a vital research direction for the future of explainable and trustworthy AI.